OpenStudy (anonymous):

2,192,536 express in Hindu-Arabic expanded form

OpenStudy (aum):

\[ 2,192,536 = \\ 2 \times 1,000,000 + 1 \times 100,000 + 9 \times 10,000 + 2 \times 1,000 + 5 \times 100 + 3 \times 10 + 6 \]